tHE NO. OF DIFFERENT SIGNALS CAN BE GIVEN USING ANY NUMBER OF FLAGS FROM 4flags of different colour is?


The answer is given as 64.


Plz.explain...

Grade:12

2 Answers

Sona Srivastava
13 years ago

signals using one flag = 4

signals using two flags = (4C2)(2!) .... 4C2-for choosing 2 flags and 2! for their arrangment

signals using three flags = (4C3)(3!)

signals using four flags = (4C4)(4!) 

adding all up gives 64  

Prasanjeet Mohanty Mohanty
13 years ago

In giving a signal, the order of the placement of colours is important,i.e a red and blue signal is different than the blue and red signal. Hence the order is very important.

Therefore the required no. of ways=4P1 +4P2 +4P3 +4P4 =4+12+24+24=64